OpenStudy (anonymous):

Math Analysis: Write an explicit formula for the nth term of the sequence. 17. 5, -2, -9, -16 18. 1, 1/3, 1/9, 1/27, 1/81

Parth (parthkohli):

17. You're adding -7 again and again, so that's the common difference. \( \color{Black}{\Rightarrow a_n = a_1 + (n - 1) d}\) Just replace a1 with the first term and d with the difference.

OpenStudy (anonymous):

so what's n? o:

OpenStudy (anonymous):

so what's n again u only told me a and d D:

OpenStudy (anonymous):

n is the number of terms.

OpenStudy (anonymous):

so n is 4?

OpenStudy (anonymous):

For number 17, yes. However, if you go farther into the sequence, the number changes, so you have to change n for whatever number in the arithmetic sequence.

OpenStudy (anonymous):

the answer says its tn=12-7n

OpenStudy (anonymous):

You have to use characteristic equations to solve.

OpenStudy (anonymous):

since it's a linear recurrence relation

OpenStudy (anonymous):

was the equation on top correct?

OpenStudy (anonymous):

yea, looks like it, but you have to use characteristic on it for explicit formula.

OpenStudy (anonymous):

i dont understand what you mean o-o

OpenStudy (anonymous):

someone said n=the number of terms

OpenStudy (anonymous):

could you show me how you did the problem

OpenStudy (anonymous):

5+(n-1)(-7) = a_n as he said. so 5-7n+7 = a_n so it s a_n = 12-7n.

OpenStudy (anonymous):

that's how you get the answer.

OpenStudy (anonymous):

if you mean getting to a_n = a_1 + (n-1)d, that's much more complicated to explain,

OpenStudy (anonymous):

the steps is: -writing out a recurrence relation, -translating it to a explicit formula,

OpenStudy (anonymous):

how about #18

OpenStudy (anonymous):

what's the answer for #18??? i got 3(1/3)^n